Jose is 3 times as old as Wally. Five years from now, The sum of their ages is 42. How old are they now?
Answer by ikleyn(52776)

.
Jose is 3 times as old as Wally. Five years from now, The sum of their ages is 42. How old are they now?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let J = Jose's age now; W = Wally's age now. Then

J = 3W,               (1)     ( "Jose is 3 times as old as Wally" )
(J+5) + (W+5) = 42.   (2)     ( "Five years from now, The sum of their ages is 42" )

Simplify 2):

J + W + 10 = 42,   or

J + W = 42-10 = 32.

Now substitute J= 3W from (1) into the last equation. You will have

3W + W = 32.

Can you complete the solution from this point?

There is a bunch of lessons with solved age problems in this site
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - OVERVIEW of lessons on age problems

They were developed to teach you on solving these problems.

Read them and become an expert in this area.